OpenStudy (anonymous):

Write the expression 125m^6 - 8n^3 in factored form.

OpenStudy (jwheele1):

\[125m ^{6}-8n ^{3}\]

OpenStudy (anonymous):

Yeah^

OpenStudy (jwheele1):

step one....is there any factors of 125 and 8 that are the same?

OpenStudy (jwheele1):

for example: if the problem was 5m+10m^2 the 5 and the 10 have a factor in common and that is 5

OpenStudy (anonymous):

I don't think there's a common factor betweem 125 and 8

OpenStudy (anonymous):

Hint: 5³ = 125 and 2³ = 8

OpenStudy (anonymous):

5 and 2?

OpenStudy (anonymous):

Yeah. Also remember that (a³ - b³) = (a-b) (a²+a b+b²)

OpenStudy (anonymous):

okay

OpenStudy (jwheele1):

so, 125 and 8 have no like factors =)

OpenStudy (jwheele1):

m is also not n....so same problem

OpenStudy (jwheele1):

it is either not factorable or she wants the factors for each term

OpenStudy (jwheele1):

5*5*5*m*m*m*m*m*m-2*2*2*n*n*n

OpenStudy (anonymous):

It's factorable...

OpenStudy (jwheele1):

for instance 5m+10m2 would have been 5m(1+5m) factored

OpenStudy (jwheele1):

how?

OpenStudy (anonymous):

Let m² = p then factor it. Hint: (a³ - b³) = (a-b) (a²+a b+b²)

OpenStudy (jwheele1):

what formula is that?

OpenStudy (jwheele1):

the hint one

OpenStudy (anonymous):

So would i do this it has nocommon factor

OpenStudy (anonymous):

@jwheele1 Perfect Cubic Polynomial http://mathworld.wolfram.com/PerfectCubicPolynomial.html

OpenStudy (anonymous):

Let m² = p \[125m^6 - 8n^3 = 125p^3 - 8n^3\] Now factor it.\[125p^3 - 8n^3 = (5p)^3 - (2n)^3 = (5p - 2n)((5p)^2 + (5p)(2n) + (2n)^2)\]

OpenStudy (anonymous):

i see.. thank you

OpenStudy (anonymous):

I know it looks complicated. Do you understand?